# ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection

### v2

@@ 4.3 Training and Inference
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel‐wise concatenation; for each layer $l$, $M_{n}^{l}$ and $M_{a}^{l}$ are the normal and anomaly score maps; $S_{\mathrm{gt}}$ is the ground‐truth mask; and $I$ is the all‐ones matrix.
During inference, the final image‐level anomaly score is computed as:$\mathrm{Score}=\frac{1}{N}\sum_{l=1}^{N}S_{a}^{l}$. For pixel‐level prediction, we merge all intermediate maps $M_{n}^{l}$ and $M_{a}^{l}$, then obtain the anomaly map $\mathrm{Map}\in\mathbb{R}^{H\times W}$ as follows:
@@ 5.1 Setup
To comprehensively evaluate ViP2-CLIP across diverse application scenarios, we conduct extensive experiments on public benchmarks spanning industrial and medical domains. In the industrial domain, we use MVTec AD [2], VisA [34], MPDD [13], BTAD [20], KSDD [26], DAGM [29], and DTD-Synthetic [1]. For medical imaging, we include brain tumour detection datasets HeadCT [25], BrainMRI [25], and Br35H [10], skin cancer detection dataset ISIC [7], colon polyp detection datasets CVC-ClinicDB [3], CVC-ColonDB [27], Endo [11], and Kvasir [14]. We compare ViP2-CLIP against five leading ZSAD methods: CLIP [23], WinCLIP [12], APRIL-GAN [5], AnomalyCLIP [33] and AdaCLIP [4]. More details about the datasets or methods are shown in Appendix B.
@@ 5.2 Main Results
Tab. 1 compares ViP2-CLIP against five representative baselines on seven industrial defect benchmarks. Our model delivers competitive results and outperforms all baselines in most datasets. The vanilla CLIP struggles, as its pre-training emphasizes general object semantics rather than anomaly patterns. WinCLIP and APRIL-GAN improve detection through handcrafted prompts and local features tuning. AnomalyCLIP employs object-agnostic prompts but remains limited by static prompt design and shallow cross-modal fusion. AdaCLIP further refines prompts in both visual and textual space, which boosts F1 but neglects AUPRO, leading to an imbalanced ZSAD. In contrast, ViP2-CLIP yields consistent improvements across all metrics, achieving state-of-the-art performance. Our ViP-Prompt creates fine-grained normal and abnormal prompts that align tightly with object features; paired with UTPA’s unified alignment, it markedly boosts accuracy at both image-level detection and pixel-level localization.
@@
Fig. 4 compares the anomaly detection results of our approach with four strong baselines, showing that ViP2-CLIP more accurately captures defects with sharper boundaries, supporting the observed quantitative gains. We also provide failure cases analyses in Appendix C.
@@
To assess cross-domain transferability, we benchmark ViP2-CLIP on eight public medical defect datasets. As shown in Tab. 2, ViP2-CLIP achieves consistently superior detection performance on most medical datasets, generally outperforming prior methods. Qualitative examples in Fig. 4 further support this claim: our model precisely localizes melanoma lesions in dermoscopy images and colonic polyps in endoscopic frames, demonstrating robust adaptability to disparate pathological patterns.
@@
To quantify the benefit of visual conditioning, we compare ViP2-CLIP against ViP2-CLIP_re, which only uses static learnable prompts without visual cues. Fig. 5 reports image-level and pixel-level F1 gains on the seven industrial datasets. ViP2-CLIP achieves higher pixel‐level F1 on all datasets and improves image‐level F1 on five of them. The slight drops on BTAD and DAGM are attributed to significant domain shifts from MVTec AD: BTAD features real-world industrial environments with complex lighting conditions, while DAGM contains synthetic textures with small, repetitive anomalies. These differences likely cause shifts in global visual embeddings, limiting ICA’s transferability and leading to imbalanced detection performance. This underscores the importance of domain consistency in transfer learning, as large domain gaps can hinder generalization. Despite minor image-level drops on a few datasets, ViP-Prompt consistently achieves SOTA performance across diverse detection scenarios, demonstrating that incorporating both global and local visual cues into prompts enhances cross-modal alignment and promotes robust generalization.
@@
Computational Efficiency Analysis
Tab. 3 compares the computational efficiency of the top three baselines. AnomalyCLIP modifies self-attention weights via DPAM, while AdaCLIP introduces additional learnable tokens and a multi-scale clustering module, both of which increase computational overhead. In contrast, ViP2‑CLIP employs only two lightweight adapters outside the frozen CLIP backbone, achieving the fastest training (0.54 h), inference (48.7 ms/image), and the lowest GPU memory usage (2.21 GB), making it well-suited for deployment in resource-constrained industrial environments.
5.3 Ablation Study & Discussion
Module Ablation
@@ 5.3 Ablation Study & Discussion
UTPA for Resolving Optimization Conflicts
We conduct an ablation study using a frozen CLIP backbone with learnable prompts for normal and anomalous classes to evaluate the effectiveness of UTPA in guiding network learning. The baseline (CLIP_DUAL) adopts a dual-branch strategy that aligns prompts with both global and local visual embeddings. In contrast, CLIP_UTPA simplifies the optimization by aligning solely with local features. As shown in Tab. 5, on the MVTec AD dataset, the dual-branch strategy tends to overfit image-level scores while underperforming on pixel-level detection. UTPA, by comparison, achieves more stable and superior results. As for VisA, it outperforms the baseline across both pixel- and image-level metrics. Several results in Tab. 4 also support this observation: traditional dual-branch approaches often suffer from imbalanced performance due to conflicting objectives, whereas UTPA alleviates this issue through consistent text-patch alignment. This facilitates more robust model learning and provides a novel alignment paradigm for more accurate ZSAD research.
Top‐K Ablation
To assess how the image-level representation is affected by the number of anomalous patches, we vary $K\in\{1,10,20,50,100,200\}$ and examine the AUROC performance in Fig. 6. At $K=1$, the model is overly sensitive to local noise, yielding a low AUROC. As $K$ increases, performance steadily improves, demonstrating that aggregating multiple high-scoring anomalous patches enhances robustness. On MVTec AD, performance plateaus when $K$ is at 50 or 100; As for VisA, the optimum occurs at $K=50$. Beyond these points, additional patches introduce mostly normal regions, diluting the anomaly signals and reducing accuracy. Balanced across all metrics, we select the top 50 most anomalous patches as image-level descriptors, which aligns well with the anomaly distribution characteristics in most detection scenarios.
Feature Visualization at Different layers
To investigate the role of network depth in capturing anomalous features, Fig. 7 presents anomaly maps derived from $F_{m}^{1}$ to $F_{m}^{4}$ layers. We observe that for complex object categories such as hazelnut, deeper layers yield more accurate anomaly localization. In contrast, for simpler texture-based objects like leather, intermediate layers offer sufficient discriminative power. Aggregating features across multiple depths provides complementary contextual cues, allowing the model to better adapt to diverse anomaly types, and thereby achieve more robust anomaly detection.
State‐Adjective Ablation
Figure 7: Visualization of anomaly maps from different layers.

### v3

@@ 4.3 Training and Inference
where $\cdot$$\cdot$] denotes channel‐wise concatenation; for each layer $l$, $M_{n}^{l}$ and $M_{a}^{l}$ are the normal and anomaly score maps; $S_{\mathrm{gt}}$ is the ground‐truth mask; and $I$ is the all‐ones matrix.
[…]
@@ 5.1 Setup
To comprehensively evaluate ViP2-CLIP across diverse application scenarios, we conduct extensive experiments on public benchmarks spanning industrial and medical domains. In the industrial domain, we use MVTec AD [2], VisA [38], MPDD [15], BTAD [23], KSDD [30], DAGM [33], and DTD-Synthetic [1]. For medical imaging, we include brain tumour detection datasets HeadCT [29], BrainMRI [29], and Br35H [12], skin cancer detection dataset ISIC [7], colon polyp detection datasets CVC-ClinicDB [3], CVC-ColonDB [31], Endo [13], and Kvasir [16]. We compare ViP2-CLIP against six leading ZSAD methods: CLIP [26], WinCLIP [14], APRIL-GAN [5], AnomalyCLIP [37], AdaCLIP [4] and AA-CLIP [22]. Additional details regarding the datasets and methodological specifics are provided in Appendix B.
@@ 5.2 Main Results
Tab. 1 compares ViP2-CLIP against six representative baselines on seven industrial defect benchmarks. Our model delivers competitive results and outperforms all baselines in most datasets. The vanilla CLIP struggles, as its pre-training emphasizes general object semantics rather than anomaly patterns. WinCLIP and APRIL-GAN improve detection through handcrafted prompts and local features tuning. AnomalyCLIP employs object-agnostic prompts but remains limited by static prompt design and shallow cross-modal fusion. AA-CLIP directly introduces adapter modules to enlarge the semantic gap between normal and anomalous texts, thereby enhancing detection performance. AdaCLIP refines prompts in both visual and textual space, which boosts F1 but neglects AUPRO, leading to an imbalanced ZSAD. In contrast, ViP2-CLIP yields consistent improvements across all metrics, achieving state-of-the-art performance. Our ViP-Prompt creates fine-grained normal and abnormal prompts that align tightly with object features; paired with UTPA’s unified alignment strategy, substantially improves overall ZSAD performance.
@@
Fig. 4 further compares the anomaly detection results of our approach with five strong baselines, showing that ViP2-CLIP more accurately captures defects with sharper boundaries, supporting the observed quantitative gains. We also provide failure-case analyses in Appendix C.
@@
To assess cross-domain transferability, we benchmark ViP2-CLIP on eight public medical defect datasets. As shown in Tab. 2, although AA-CLIP achieves stronger pixel-level anomaly localization, its weaker image-level detection results in unbalanced ZSAD performance. In contrast, our ViP2-CLIP consistently delivers superior results across most datasets. Qualitative examples in Fig. 4 also support this claim: our model precisely localizes melanoma lesions in dermoscopy images and colonic polyps in endoscopic frames, demonstrating robust adaptability to disparate pathological patterns.
@@
To quantify the benefit of visual conditioning, we compare ViP2-CLIP against ViP2-CLIP_re, which only uses static learnable prompts without visual cues. Fig. 5 reports image-level and pixel-level F1 gains on the seven industrial datasets. ViP2-CLIP achieves higher pixel‐level F1 on all datasets and improves image‐level F1 on five of them. The slight drops on BTAD and DAGM are attributed to significant domain shifts from MVTec AD: BTAD reflects real-world industrial settings with complex lighting conditions, while DAGM contains synthetic textures characterized by small, repetitive anomalies. Such shifts likely distort global visual embeddings, reducing the transferability of ICA and resulting in imbalanced performance. Despite minor image-level declines on a few datasets, ViP-Prompt consistently achieves SOTA performance across diverse detection scenarios. It effectively mitigates segmentation fluctuations arising from uncertain class names in prompts under data privacy constraints (as detailed in Appendix A), thereby ensuring more robust generalization.
Computational Efficiency Analysis
Tab. 3 compares the computational efficiency with the top three baselines. AnomalyCLIP modifies self-attention weights via DPAM, while AdaCLIP introduces additional learnable tokens and a multi-scale clustering module, both of which increase computational overhead. Although AA-CLIP attains the shortest inference time through a simplified network design, its large number of trainable parameters diminishes overall efficiency. In contrast, ViP2‑CLIP employs only two lightweight adapters outside the frozen CLIP backbone, enabling the fastest training (0.54 h), competitive inference speed (48.7 ms/image), and the lowest GPU memory consumption (2.21 GB), making it highly suitable for deployment in resource-constrained industrial environments.
@@ 5.3 Ablation Study & Discussion
UTPA for Resolving Optimization Conflicts
We conduct an ablation study using a frozen CLIP backbone with learnable prompts for normal and anomalous classes to evaluate the effectiveness of UTPA in guiding model learning. The baseline (CLIP_DUAL) adopts a dual-branch strategy that aligns prompts with both global and local visual cues. In contrast, CLIP_UTPA simplifies the optimization by aligning solely with local features. As shown in Tab. 13, on the MVTec AD dataset, the dual-branch strategy tends to overfit image-level scores while underperforming on pixel-level detection. UTPA, by comparison, achieves more stable and superior results. As for VisA, it outperforms the baseline across both detection metrics. Several results in Tab. 4 also support this observation: traditional dual-branch approaches often suffer from imbalanced performance due to conflicting objectives, whereas UTPA alleviates this issue through consistent text-patch alignment. This facilitates more robust model learning and provides a novel alignment paradigm for more accurate ZSAD research.
Top‐K Ablation
To assess how the image-level representation is affected by the number of anomalous patches, we vary $K\in\{1,10,20,50,100,200\}$ and examine the AUROC performance in Fig. 6. At $K=1$, the model is overly sensitive to local noise, yielding a low AUROC. As $K$ increases, performance steadily improves, demonstrating that aggregating multiple high-scoring anomalous patches enhances robustness. On MVTec AD, performance plateaus when $K$ is at 50 or 100; As for VisA, the optimum occurs at $K=50$. Beyond these points, additional patches introduce mostly normal regions, diluting the anomaly signals and reducing accuracy. Balanced across all metrics, we select the top 50 most anomalous patches as image-level descriptors, which effectively capture the anomaly characteristics in most detection scenarios.
Visualization at Different layers
To examine the impact of model depth on anomaly representation, Fig. 7 presents anomaly maps derived from $F_{m}^{1}$ to $F_{m}^{4}$ layers. For complex object categories such as hazelnut, deeper layers yield more accurate anomaly localization. In contrast, for simpler texture-based objects like leather, intermediate layers offer sufficient discriminative power. Aggregating features across multiple depths provides complementary contextual cues, allowing the model to better adapt to diverse anomaly types, and thereby achieve more robust anomaly detection.
Figure 7: Visualization of anomaly maps from different layers.
@@
6 Conclusion
We propose ViP2‑CLIP, a universal ZSAD framework that detects anomalies in unseen categories without any target domain training samples. At its core, ViP‑Prompt adaptively integrates global and local visual cues into text prompts, eliminating reliance on manual templates or class-name priors, achieving superior generalization and robustness in cross-modal alignment. In addition, the UTPA optimizes a unified text-patch alignment strategy across scales, sharpening both detection and localization performance. Extensive experiments on 15 mainstream datasets demonstrate that ViP2‑CLIP consistently outperforms existing methods across various detection scenarios, especially when category labels are ambiguous or privacy-constrained, offering a more accurate and scalable solution for ZSAD.
Limitations
Although UTPA achieves high-precision ZSAD performance through a simple yet effective alignment strategy, its current implementation only selects the better-performing Top-K from a small set of discrete candidates, further work on more advanced feature aggregation strategies is needed to explore.
Appendix A Motivation Statement
A.1 Robustness Analysis of Class-Name in Prompts
Methods like CLIP-AD [6], AdaCLIP [4], and KANOCLIP [19] enhance alignment by embedding class names into prompts, but this introduces a critical vulnerability: CLIP’s segmentation quality is highly sensitive to label phrasing. Specifically, we conducted synonym-swap experiments on the training-free WinCLIP model to evaluate how CLIP’s inherent segmentation capability is influenced by the class-name variations. As shown in Fig. 8, by replacing class names with semantically equivalent variants (e.g., ‘zipper’ to ‘zip’ or ‘fastener’), we observed a striking inconsistency: pixel-level segmentation metrics can vary by up to 10 percentage points, while image-level classification metrics remain stable. We attribute this instability to CLIP’s classification-oriented pre-training, which enforces robust global alignment but yields fragile local alignment.
Figure 8: WinCLIP’s performance of different Class Names in Prompts.
To further evaluate whether training-based modality alignment can alleviate this sensitivity, we repeated the synonym-swap experiments on recent work, AdaCLIP. As shown in Fig. 9, although AdaCLIP’s pixel-level F1 stays stable, its AUROC and PRO fluctuate with label swaps; in particular, changing ‘carpet’ to ‘floor mat’ triggers a 47% swing in PRO. This dramatic fluctuation shows that, even with fine-tuning on auxiliary data, the limited scale of anomaly-specific training cannot effectively mitigate CLIP’s segmentation sensitivity to class-name perturbations. These findings underscore the necessity of label-agnostic prompt designs for truly robust ZSAD.
Figure 9: AdaCLIP’s performance of different Class Names in Prompts.
Appendix B Experimental Details
Domain Dataset Category Modalities $|\mathcal{C}|$ Normal and anomalous samples
Industrial MVTec AD Obj & texture Photography 15 (467, 1258)
VisA Obj Photography 12 (962, 1200)
MPDD Photography 6 (176, 282)
BTAD Photography 3 (451, 290)
KSDD Photography 1 (286, 54)
DAGM Texture Photography 10 (6996, 1054)
DTD-Synthetic Photography 12 (357, 947)
Medical ISIC Skin Photography 1 (0, 379)
CVC-ClinicDB Colon Endoscopy 1 (0, 612)
CVC-ColonDB Endoscopy 1 (0, 380)
Kvasir Endoscopy 1 (0, 1000)
Endo Endoscopy 1 (0, 200)
HeadCT Brain Radiology (CT) 1 (100, 100)
BrainMRI Radiology (MRI) 1 (98, 155)
Br35H Radiology (MRI) 1 (1500, 1500)
Table 6: Statistics of the datasets.
B.1 Implementation Details
We use the publicly released CLIP (ViT‑L/14@336px) model as the frozen feature extractor. In all experiments, the learnable prompt length is set to 10, of which 3 tokens fuse with the global visual feature. In UTPA, we set Topk=50Top‑k=50 and perform cross-modal alignment at layers $\{6,12,18,24\}$. During training, input images are resized to $518\times 518$, we optimize with Adam at a learning rate of $1\times 10^{-3}$ for 10 epochs with a batch size of 8. All experiments are implemented in PyTorch 2.6.0 on a single NVIDIA L20 (48 GB).
We conduct comprehensive evaluations across 15 publicly available benchmarks spanning industrial and medical domains. Specifically, we fine‑tune on MVTec AD’s test set and perform zero‑shot evaluation on all other datasets; for MVTec AD, we fine‑tune on VisA’s test split. We adopt standard metrics for evaluation: Area Under the ROC Curve (AUC), Average Precision (AP), and F1-score (F1) for image-level anomaly detection; AUC, Per-Region Overlap (PRO), and F1 for pixel-level segmentation.
B.2 Baselines
To validate the superiority of ViP2‑CLIP, we compare against the following state‑of‑the‑art methods:
  • CLIP [26]: We adapt CLIP for anomaly detection by using two prompt templates—‘a photo of a normal [cls]’ and ‘a photo of an abnormal [cls]’, where [cls] is the category class-name. We compute image‑level anomaly scores via cosine similarity between global visual features and text embeddings. For pixel‑level segmentation, we extend this scoring to the last‑layer patch embeddings.
  • WinCLIP [14]: WinCLIP is the first CLIP‑based ZSAD method, which uses multiple manually designed prompts and a multi‑window feature extraction strategy. All parameters are kept the same as in their paper.
  • APRIL‑GAN[5]: APRIL‑GAN extends WinCLIP by incorporating learnable linear projection layers to enrich local visual feature representations. All experiments are conducted using the official weights and adhere strictly to the original implementation protocols.
  • AnomalyCLIP [37]: AnomalyCLIP introduces object‑agnostic learnable prompts and a DPAM mechanism for stronger local features’ modelling. All experiments are conducted using publicly released official model weights.
  • AdaCLIP [4]: AdaCLIP incorporates hybrid prompts into both text and image encoders with an HSF module for improved cross‑modal fusion. As AdaCLIP adopts an evaluation protocol distinct from ours, it trains on MVTec AD and CVC-ColonDB and performs zero-shot testing on the remaining datasets. To ensure fair and consistent comparison, we reimplement AdaCLIP under our unified evaluation protocol. All parameters are kept the same as in their paper.
  • AA-CLIP [22]: AA-CLIP enhances anomaly detection by adopting a two-stage training strategy that enlarges the semantic gap between normal and anomalous texts through joint optimization of the text and image encoders. We reimplement it under our unified evaluation protocol, keeping all parameters consistent with the original paper.
Task Category Datasets PatchCore RD4AD ViP2-CLIP
Image-level
(AUROC, AP)
Obj &texture MVTec AD (99.0, 99.7) (98.7, 99.4) (91.2, 96.0)
Obj VisA (94.6, 95.9) (95.3, 95.7) (88.5, 90.4)
MPDD (94.1, 96.3) (91.6, 93.8) (79.7, 84.5)
BTAD (93.2, 98.6) (93.8, 96.8) (95.0, 98.4)
Texture DAGM (92.7, 81.3) (92.9, 79.1) (98.5, 94.3)
Pixel-level
(AUROC, PRO)
Obj &texture MVTec AD (98.1, 92.8) (97.8, 93.6) (90.5, 87.1)
Obj VisA (98.5, 92.2) (98.4, 91.2) (95.4, 92.2)
MPDD (98.8, 94.9) (98.4, 95.2) (97.2, 92.6)
BTAD (97.4, 74.4) (97.5, 75.1) (95.6, 86.1)
Texture DAGM (95.9, 87.9) (96.8, 91.9) (97.5, 95.2)
Table 7: Comparison of ZSAD performance between ViP2-CLIP and SOTA full-shot methods.Results of methods with * are copied from the AnomalyCLIP paper.The best performance is highlighted in red, and the second is highlighted in blue.
B.3 Datasets
We evaluate on 7 industrial datasets (MVTec AD, VisA, MPDD, BTAD, KSDD, DAGM, DTD‑Synthetic) and 8 medical datasets (HeadCT, BrainMRI, Br35H, ISIC, CVC‑ClinicDB, CVC‑ColonDB, Endo, Kvasir), for a total of 15 benchmarks. All methods are trained and tested solely on each dataset’s test split. The dataset statistics are given in Tab. 6. We apply OpenCLIP’s default normalization to all images and resize them to $(518,518)$ to obtain suitably scaled feature maps.
Appendix C Additional Experiments and Analysis
C.1 Comparison with Full-shot Methods
In this section, we compare ViP2‑CLIP against two state‑of‑the‑art full-shot methods, PatchCore [28] and RD4AD [8], on five industrial datasets for which training samples are available. Results are presented in Tab. 7. Despite not using any target domain training samples, ViP2‑CLIP achieves detection and segmentation performance on par with or superior to these fully supervised approaches, with particularly strong gains on the BTAD and DAGM datasets. This demonstrates that ViP2‑CLIP’s visual‑perception prompt mechanism adaptively generates fine‑grained normal and anomalous text descriptions that fully unleash CLIP’s ZSAD performance, breaking the dependency of traditional methods on in‑domain training data and exhibits excellent generalization.
Task Datasets $|\mathcal{C}|$ AnomalyCLIP ViP2-CLIP
Image-level
(AUROC, AP, F1)
HeadCT 1 (94.3, 94.0, 90.5) (96.5, 96.5, 92.8)
BrainMRI 1 (94.7, 95.7, 92.5) (96.7, 97.6, 94.9)
Brain35H 1 (96.5, 96.7, 92.6) (97.1, 97.5, 93.4)
AVERAGE - (95.2, 95.5, 91.9) (96.8, 97.2, 93.7)
Pixel-level
(AUROC, PRO, F1)
ISIC 1 (87.8, 75.6, 69.7) (91.7, 82.9, 75.3)
CVC-ColonDB 1 (87.8, 76.9, 48.5) (87.7, 86.2, 53.5)
CVC-ClinicDB 1 (88.8, 76.1, 51.7) (92.2, 84.3, 60.8)
Endo 1 (89.0, 73.5, 58.1) (92.3, 82.8, 65.2)
Kvasir 1 (85.6, 50.0, 55.1) (90.5, 56.9, 63.4)
AVERAGE - (87.8, 70.4, 56.6) (90.9, 78.6, 63.6)
Table 8: ZSAD performance on medical images after fine-tuning by medical-domain datasets. Best performance is shown in red.
C.2 Targeted Fine‑tuning on Medical Data
Although ViP2‑CLIP achieves strong results on industrial benchmarks, its performance on medical images is relatively weaker, partly due to the domain gap between industrial data and medical target domains. To investigate whether fine‑tuning on medical auxiliary data can bridge this gap, we design a cross‑domain fine-tuning protocol: we train on MVTec AD and CVC-ColonDB test split and test on the remaining medical datasets; for CVC-ColonDB, we use VisA and CVC-ClinicDB test sets as auxiliary data. We compare against AnomalyCLIP under the same training strategy. As shown in Tab. 8, introducing medical auxiliary data significantly boosts detection performance, particularly on CVC-ClinicDB and Endo datasets, highlighting the critical role of auxiliary data in cross‑domain generalization. Overall, ViP2‑CLIP outperforms AnomalyCLIP under medical fine‑tuning, demonstrating superior cross‑domain anomaly detection capability.
Figure 10: Prompt Length ablation.